# When Robustness Meets Conservativeness: Conformalized Uncertainty Calibration for Balanced Decision Making

### v1

@@ 5 Conclusion
We proposed CREME, a conformalized framework for robust decision-making that certifies both miscoverage and regret across robustness levels, enabling practitioners to select appropriate robustness in optimization design. Theoretically, we showed that under mild conditions, the estimator provides conservative, distribution-free finite-sample guarantees and traces a certified Pareto frontier for the miscoverage–regret tradeoff. Empirically, across four canonical optimization paradigms, CREME produces frontiers with high validity and accuracy and guides robustness selection to balance miscoverage and regret. Thus, CREME offers a principled alternative to ad hoc calibration, addressing a key gap in the existing literature.
We note that while CREME itself is a lightweight procedure, its computational efficiency is ultimately bottlenecked by repeatedly solving the underlying optimization problem and its robust variant. Specifically, for each robustness level $\lambda$, the estimator requires solving up to $2n$ optimization problems, where $n$ is the number of calibration samples. This cost can become substantial in high-dimensional or nonconvex settings. Nevertheless, in our experiments, spanning a wide range of convex optimization tasks, we find that CREME remains practical with standard solvers, and becomes particularly efficient when closed-form regret solutions are available. Looking ahead, we expect future work to focus on enhancing the scalability of CREME, thereby extending its applicability to more complex and large-scale decision problems.
Supplementary Materials for
“Conformalized Uncertainty Calibration
for Balanced Decision Making”
Appendix A Illustrative Examples: When the Assumptions Hold (and When They Fail)
Example (assumption holds)
Consider a linear optimization with a compact convex feasible region
$\min_{z}~\langle Y,z\rangle\quad\text{s.t.}\quad\|z\|_{2}\leq M,$
and assume $\|Y\|\leq K$ a.s. (hence regret is uniformly bounded) and that measurable selectors are taken. We define nested uncertainty sets $\mathcal{C}_{\lambda}=\{y:\|y-\mu\|_{2}\leq\lambda\}$, which by definition guarantees that the miscoverage is a decreasing function of $\lambda$. We define the robust variant of the optimization problem as
$\min_{z}\max_{y\in\mathcal{C}_{\lambda}}~\langle y,z\rangle\quad\text{s.t.}\quad\|z\|_{2}\leq M.$
@@ Appendix A Illustrative Examples: When the Assumptions Hold (and When They Fail)
Therefore, $\mathbb{E}[f(Y,z_{\lambda}^{*})]=\langle\mu,z_{\lambda}^{*}\rangle\in\{0,-M\|\mu\|_{2}\}$ is weakly increasing as $\lambda$ increasing, satisfying 3. Nesting holds by construction, and
so $R_{\lambda}(X,Y)\leq B$ by setting $B=2MK$, and it is easy to verify that $R_{\lambda}(X,Y)\leq B$ is an increasing function of $\lambda$, therefore the statements in Proposition 3 hold. In fact, the frontier in this example is strictly decreasing whenever $\mathbb{P}\{\|\mu\|_{2}>\lambda\}>0$ for some $\lambda$. This example illustrates that the assumption is mild and naturally satisfied in common robust linear programs with ellipsoidal uncertainty and bounded outcomes.
Remark 1.
Counterexample (assumption fails)
Define $\Lambda=\{\lambda_{1},\lambda_{2}\}$ with $\lambda_{1}<\lambda_{2}$, and nested uncertainty sets $\mathcal{C}_{\lambda_{1}}=\{-1\}$ and $\mathcal{C}_{\lambda_{2}}=\{-1,+1\}$. It can be seen that the (robust) solutions to $\min_{z}\max_{y\in\mathcal{C}_{\lambda}}f(y,z)$ can be divided into two cases: ($i$) When $\lambda=\lambda_{1}$, $z^{*}_{\lambda_{1}}=-1$; ($ii$) When $\lambda=\lambda_{2}$, $z^{*}_{\lambda_{2}}=0$. Therefore, it can be seen that
which makes this optimization problem violate 3.
Appendix B Proof for Proposition 2
Proof for Proposition 2.
@@ Appendix B Proof for Proposition 2
where for the last inequality we used the bounding condition (2). Observe that by the i.i.d. condition, using Hoeffding’s inequality, we know that for any $\delta>0$, there is
@@
Using Proposition 2, we can further establish a reliable lower bound on the difference in the true expected loss function between two robustness parameters, as summarized below.
Corollary 2.
@@
Proof of Corollary 2.
Since for any $\lambda\in\Lambda$, we know from Proposition 2 that with probability $1-\delta^{\prime}$, there is
@@
The difference in the true expected loss function between two robustness parameters is particularly useful in practice, as it reflects how much gain or loss a decision-maker can expect when switching from $\lambda_{1}$ to $\lambda_{2}$, as illustrated in Figure 1. Corollary 2 suggests that we can simply use the difference in estimators after offsetting with the error term $2\epsilon$ to obtain a lower bound, which informs the decision maker of a conservative estimate of the change magnitude. Practically, this can help reliably assess the trade-offs involved in selecting different robustness levels, leading to more principled and robust decision-making.
Appendix C Proof for Certified Pareto Frontier
Proof for Proposition 3.
By nesting, for any $(X,Y)$ and $\lambda_{1}\leq\lambda_{2}$, $\mathcal{C}_{\lambda_{1}}(X)\subseteq\mathcal{C}_{\lambda_{2}}(X)$, hence
$\mathbbm{1}\{Y\notin\mathcal{C}_{\lambda_{1}}(X)\}\geq\mathbbm{1}\{Y\notin\mathcal{C}_{\lambda_{2}}(X)\}.$
@@ Appendix C Proof for Certified Pareto Frontier
The second term is independent of $\lambda$. By the law of iterated expectations and Assumption 3,
@@
Because as $\lambda$ increases (i.e., $\mathcal{C}_{\lambda}$ expands), $\alpha_{I}$ weakly decreases while $\alpha_{R}$ weakly increases, the map $\lambda\mapsto(\alpha_{I}(\lambda),\alpha_{R}(\lambda))$ is (weakly) decreasing. Fix $\lambda\in\Lambda$. For any $\lambda^{\prime}\neq\lambda$, if $\lambda^{\prime}<\lambda$ then $\alpha_{I}(\lambda^{\prime})\geq\alpha_{I}(\lambda)$ and $\alpha_{R}(\lambda^{\prime})\leq\alpha_{R}(\lambda)$; if $\lambda^{\prime}>\lambda$ the inequalities reverse. Hence no policy in the family $\{z^{*}_{\lambda}\}$ strictly improves both coordinates over $z^{*}_{\lambda}$. Therefore every image point is Pareto efficient, and the image is exactly the Pareto frontier achievable by this policy family. (Ties may occur if the inequalities become equalities for some $\lambda_{1}<\lambda_{2}$.) ∎
Proof for Corollary 1.
@@
Appendix D Additional Experiment Details
@@
Similar to Appendix A, the synthetic data used in our experiments focuses on settings where $X$ is degenerated, so we only have observations of $Y$. As we’ve mentioned before, this simplification does not limit the generality of our experimental findings, since both the Pareto frontier and its associated performance metrics are inherently unconditioned on $X$. The implementation is developed in Python, with core dependencies on the cvxpy [Diamond and Boyd, 2016] and cvxpylayers [Agrawal et al., 2019] libraries, which are used to solve convex optimization problems. We use the default parameter settings in their libraries across all experiments. Across all settings, the upper bound $B$ is heuristically set to the maximum observed regret for a large number ($1\times 10^{2}$) of simulations.
D.1 Experimental Setups
@@
In this section, we present the detailed experimental setups for the four optimization problems considered in our study. Across all settings, we specify the uncertainty set as $\mathcal{C}_{\lambda}=\{\tilde{y}:\|\tilde{y}-\mu\|_{\infty}\leq\lambda\}$, where $\mu$ denotes the mean of the distribution of the outcome variable $Y$. These optimizations are solved via convex solvers in our implementation.
@@
$\min_{z\in\mathbb{R}^{2}}\max_{y\in\mathcal{C}_{\lambda}}\;y^{\top}z\quad\text{s.t.}\quad Az\leq b.$
@@
$\min_{z\in\mathbb{R}_{+}}\;\max_{y\in\mathcal{C}_{\lambda}}\;\big[-p\min(y,z)+cz-v(z-y)^{+}\big].$
@@
$\min_{z\in\mathbb{R}_{+}^{m}}\;\max_{y\in\mathcal{C}_{\lambda}}\;y^{\top}z\quad\text{s.t.}\quad Az=b.\quad\Longleftrightarrow\quad\min_{z\in\mathbb{R}_{+}^{n_{\text{edge}}}}\;\mu^{\top}z+\lambda\,(w^{\top}z)\quad\text{s.t.}\quad Az=b,$
@@
D.2 Additional Experiment Results

### v2

@@ 5 Conclusion
We proposed CREME, a conformalized framework for robust decision-making that certifies both miscoverage and regret across robustness levels, enabling practitioners to select appropriate robustness in optimization design. Theoretically, we showed that under mild conditions, the estimator provides conservative, distribution-free finite-sample guarantees and traces a certified Pareto frontier for the miscoverage–regret tradeoff. Across four optimization implementations, CREME produces frontiers with high validity and accuracy and guides robustness selection to balance miscoverage and regret. Thus, CREME offers a principled alternative to ad hoc calibration, addressing a key gap in the existing literature.
A limitation of this work is its reliance on the essential supremum bound on the regret (3.2), an assumption inherited from the original conformal risk control framework (Angelopoulos et al., 2022). An important direction for future research is relaxing this assumption or developing alternative conformalized procedures that achieve comparably conservative guarantees without requiring this bound.
Appendix A Proof of Theorem 3.4
Proof.
We prove by showing:
$\mathbb{E}\left[\hat{\alpha}_{\ell}(\lambda)\right]\overset{(a)}{\geq}\mathbb{E}\left[\tilde{\alpha}_{\ell}(\lambda)\right]\overset{(b)}{\geq}\mathbb{E}[\ell_{\lambda}(X_{n+1},Y_{n+1})].$ (11)
@@
therefore $\hat{\alpha}_{\ell}(\lambda)\geq\tilde{\alpha}_{\ell}(\lambda)$ almost surely, which implies $(a)$. To prove $(b)$, notice that 3.1 implies
$\mathbb{E}\left[\ell_{\lambda}(X_{1},Y_{1})\right]=\ldots=\mathbb{E}\left[\ell_{\lambda}(X_{n},Y_{n})\right]=\mathbb{E}\left[\ell_{\lambda}(X_{n+1},Y_{n+1})\right],$ (12)
we then expand the difference of two sides of $(b)$ by using the definition of $\tilde{\alpha}_{\ell}(\lambda)$ in (6):
$\mathbb{E}\left[\tilde{\alpha}_{\ell}(\lambda)\right]-\mathbb{E}[\ell_{\lambda}(X_{n+1},Y_{n+1})]=\frac{1}{n+1}\sum_{i=1}^{n}\mathbb{E}[\ell_{\lambda}(X_{i},Y_{i})]+\frac{B}{n+1}-\mathbb{E}[\ell_{\lambda}(X_{n+1},Y_{n+1})]$
$=$ $~\frac{n}{n+1}\mathbb{E}[\ell_{\lambda}(X_{n+1},Y_{n+1})]+\frac{B}{n+1}-\mathbb{E}[\ell_{\lambda}(X_{n+1},Y_{n+1})]=\frac{B-\mathbb{E}[\ell_{\lambda}(X_{n+1},Y_{n+1})]}{n+1}\geq 0.$
where the last inequality follows from 3.2. This proves $(b)$ holds and thus concludes the proof. ∎
Appendix B Proof of Proposition 3.5
Proof of Proposition 3.5.
@@
where for the last inequality we used the bounding condition (3.2). Observe that by the i.i.d. condition, using Hoeffding’s inequality, we know that for any $\delta>0$, there is
@@
Using Proposition 3.5, we can further establish a reliable lower bound on the difference in the true expected loss function between two robustness parameters, as summarized below.
Corollary B.1.
@@
Proof of Corollary B.1.
Since for any $\lambda\in\Lambda$, we know from Proposition 3.5 that with probability $1-\delta^{\prime}$, there is
@@
The difference in the true expected loss function between two robustness parameters is particularly useful in practice, as it reflects how much gain or loss a decision-maker can expect when switching from $\lambda_{1}$ to $\lambda_{2}$, as illustrated in Figure 1. Corollary B.1 suggests that we can simply use the difference in estimators after offsetting with the error term $2\epsilon$ to obtain a lower bound, which informs the decision maker of a conservative estimate of the change magnitude. Practically, this can help reliably assess the trade-offs involved in selecting different robustness levels, leading to more principled and robust decision-making.
Appendix C Discussion of 3.6
Example (assumption holds)
Consider a linear optimization with a compact convex feasible region
$\min_{z}~\langle Y,z\rangle\quad\text{s.t.}\quad\|z\|_{2}\leq M,$
and assume $\|Y\|\leq K$ a.s. (hence regret is uniformly bounded) and that measurable selectors are taken. We define nested uncertainty sets $\mathcal{U}_{\lambda}=\{y:\|y-\mu\|_{2}\leq\lambda\}$, which by definition guarantees that the miscoverage is a decreasing function of $\lambda$. We define the robust variant of the optimization problem as
$\min_{z}\max_{y\in\mathcal{U}_{\lambda}}~\langle y,z\rangle\quad\text{s.t.}\quad\|z\|_{2}\leq M.$
[…]
Therefore, $\mathbb{E}[f(Y,z_{\lambda}^{*})]=\langle\mu,z_{\lambda}^{*}\rangle\in\{0,-M\|\mu\|_{2}\}$ is weakly increasing as $\lambda$ increasing, satisfying 3.6. Nesting holds by construction, and
so $R_{\lambda}(X,Y)\leq B$ by setting $B=2MK$, and it is easy to verify that $R_{\lambda}(X,Y)\leq B$ is an increasing function of $\lambda$, therefore the statements in Proposition 3.7 hold. In fact, the frontier in this example is strictly decreasing whenever $\mathbb{P}\{\|\mu\|_{2}>\lambda\}>0$ for some $\lambda$. This example illustrates that the assumption is mild and naturally satisfied in common robust linear programs with ellipsoidal uncertainty and bounded outcomes.
Remark C.1.
Counterexample (assumption fails)
Define $\Lambda=\{\lambda_{1},\lambda_{2}\}$ with $\lambda_{1}<\lambda_{2}$, and nested uncertainty sets $\mathcal{U}_{\lambda_{1}}=\{-1\}$ and $\mathcal{U}_{\lambda_{2}}=\{-1,+1\}$. It can be seen that the (robust) solutions to $\min_{z}\max_{y\in\mathcal{U}_{\lambda}}f(y,z)$ can be divided into two cases: ($i$) When $\lambda=\lambda_{1}$, $z^{*}_{\lambda_{1}}=-1$; ($ii$) When $\lambda=\lambda_{2}$, $z^{*}_{\lambda_{2}}=0$. Therefore, it can be seen that
which makes this optimization problem violate 3.6.
Appendix D Proof of Certified Pareto Frontier
Proof of Proposition 3.7.
By nesting, for any $(X,Y)$ and $\lambda_{1}\leq\lambda_{2}$, $\mathcal{U}_{\lambda_{1}}(X)\subseteq\mathcal{U}_{\lambda_{2}}(X)$, hence
$\mathbbm{1}\{Y\notin\mathcal{U}_{\lambda_{1}}(X)\}\geq\mathbbm{1}\{Y\notin\mathcal{U}_{\lambda_{2}}(X)\}.$
@@ Appendix D Proof of Certified Pareto Frontier
The second term is independent of $\lambda$. By the law of iterated expectations and Assumption 3.6,
@@
Because as $\lambda$ increases (i.e., $\mathcal{U}_{\lambda}$ expands), $\alpha_{I}$ weakly decreases while $\alpha_{R}$ weakly increases, the map $\lambda\mapsto(\alpha_{I}(\lambda),\alpha_{R}(\lambda))$ is (weakly) decreasing. Fix $\lambda\in\Lambda$. For any $\lambda^{\prime}\neq\lambda$, if $\lambda^{\prime}<\lambda$ then $\alpha_{I}(\lambda^{\prime})\geq\alpha_{I}(\lambda)$ and $\alpha_{R}(\lambda^{\prime})\leq\alpha_{R}(\lambda)$; if $\lambda^{\prime}>\lambda$ the inequalities reverse. Hence no policy in the family $\{z^{*}_{\lambda}\}$ strictly improves both coordinates over $z^{*}_{\lambda}$. Therefore every image point is Pareto efficient, and the image is exactly the Pareto frontier achievable by this policy family. (Ties may occur if the inequalities become equalities for some $\lambda_{1}<\lambda_{2}$.) ∎
Proof of Corollary 3.8.
@@
Appendix E Proof of Corollary 3.9
Proof of Corollary 3.9.
By how $\hat{\lambda}$ is constructed, it is a random variable measurable with respect to the sigma field generated by $\left\{(X_{i},Y_{i})\right\}_{i=1}^{n}$. Therefore, $\left\{\ell_{\hat{\lambda}}(X_{i},Y_{i})\right\}_{i=1}^{n+1}\mid\hat{\lambda}$ is not an exchangeable sequence, which breaks (12). By using the definition of $\tilde{\alpha}_{\ell}(\lambda)$ in (6), there is:
$\mathbb{E}\left[\tilde{\alpha}_{\ell}(\hat{\lambda})\right]-\mathbb{E}[\ell_{\hat{\lambda}}(X_{n+1},Y_{n+1})]$
$=$ $~\frac{1}{n+1}\sum_{i=1}^{n}\mathbb{E}[\ell_{\hat{\lambda}}(X_{i},Y_{i})]+\frac{B}{n+1}-\mathbb{E}[\ell_{\hat{\lambda}}(X_{n+1},Y_{n+1})]$
$=$ $~\frac{1}{n+1}\sum_{i=1}^{n}\mathbb{E}[\ell_{\hat{\lambda}}(X_{i},Y_{i})]+\frac{B}{n+1}-\frac{1}{n+1}\sum_{i=1}^{n+1}\mathbb{E}[\ell_{\hat{\lambda}}(X_{n+1},Y_{n+1})]$
$=$ $~\frac{B-\mathbb{E}[\ell_{\hat{\lambda}}(X_{n+1},Y_{n+1})]}{n+1}+\frac{1}{n+1}\sum_{i=1}^{n}\left\{\mathbb{E}[\ell_{\hat{\lambda}}(X_{i},Y_{i})]-\mathbb{E}[\ell_{\hat{\lambda}}(X_{n+1},Y_{n+1})]\right\}$
$\geq$ $~-\frac{1}{n+1}\sum_{i=1}^{n}\left|\mathbb{E}[\ell_{\hat{\lambda}}(X_{i},Y_{i})]-\mathbb{E}[\ell_{\hat{\lambda}}(X_{n+1},Y_{n+1})]\right|$
$\geq$ $~-\frac{n}{n+1}\sup_{1\leq i\leq n}\left|\mathbb{E}[\ell_{\hat{\lambda}}(X_{i},Y_{i})]-\mathbb{E}[\ell_{\hat{\lambda}}(X_{n+1},Y_{n+1})]\right|$
$\geq$ $~-\sup_{1\leq i\leq n}\left|\mathbb{E}[\ell_{\hat{\lambda}}(X_{i},Y_{i})]-\mathbb{E}[\ell_{\hat{\lambda}}(X_{n+1},Y_{n+1})]\right|.$
The first inequality follows from 3.2. Therefore, using the chain equality (11), we conclude that
$\mathbb{E}\left[\hat{\alpha}_{\ell}(\lambda)\right]\geq\mathbb{E}[\ell_{\hat{\lambda}}(X_{n+1},Y_{n+1})]-\sup_{1\leq i\leq n}\left|\mathbb{E}[\ell_{\hat{\lambda}}(X_{i},Y_{i})]-\mathbb{E}[\ell_{\hat{\lambda}}(X_{n+1},Y_{n+1})]\right|.$
This finishes the proof. ∎
Remark E.1.
The splitting procedure (Algorithm 1) can retain the exact validity guarantee as in Theorem 3.4 because conditional exchangeability is preserved, and therefore (12) holds. Specifically, given two random splits from the calibration data, denoted by the index sets $\mathcal{I}_{1}$ and $\mathcal{I}_{2}$, then by construction of $\hat{\alpha}_{\ell}^{(1)}(\lambda)$ and $\hat{\mathcal{F}}^{(1)}$, $\hat{\lambda}$ is a random variable that is measurable with respect to the sigma field generated by $\left\{(X_{i},Y_{i})\right\}_{i\in\mathcal{I}_{1}}$. Therefore, there is
$\left\{\ell_{\hat{\lambda}}(X_{i},Y_{i})\right\}_{i\in\mathcal{I}_{2}}\mid\left\{(X_{i},Y_{i})\right\}_{i\in\mathcal{I}_{1}}\quad\text{is exchangeable}\Longrightarrow\left\{\ell_{\hat{\lambda}}(X_{i},Y_{i})\right\}_{i\in\mathcal{I}_{2}}\mid\hat{\lambda}\quad\text{is exchangeable}.$
This implies the following equality chain
$\mathbb{E}\left[\ell_{\hat{\lambda}}(X_{i_{1}},Y_{i_{1}})\mid\hat{\lambda}\right]=\ldots==\mathbb{E}\left[\ell_{\hat{\lambda}}(X_{i_{|\mathcal{I}_{2}|}},Y_{i_{|\mathcal{I}_{2}|}})\mid\hat{\lambda}\right$],
Therefore, conditioning on any $\hat{\lambda}$, there is
$\mathbb{E}[\tilde{\alpha}_{\ell}(\hat{\lambda})\mid\hat{\lambda}]-\mathbb{E}[\ell_{\hat{\lambda}}(X_{n+1},Y_{n+1})\mid\hat{\lambda}]=\frac{1}{|\mathcal{I}_{2}|+1}\sum_{i\in\mathcal{I}_{2}}\mathbb{E}[\ell_{{\hat{\lambda}}}(X_{i},Y_{i})\mid\hat{\lambda}]+\frac{B}{|\mathcal{I}_{2}|+1}-\mathbb{E}[\ell_{\hat{\lambda}}(X_{n+1},Y_{n+1})\mid\hat{\lambda}$]
$=$ $~\frac{|\mathcal{I}_{2}|}{|\mathcal{I}_{2}|+1}\mathbb{E}[\ell_{\hat{\lambda}}(X_{n+1},Y_{n+1})\mid\hat{\lambda}]+\frac{B}{|\mathcal{I}_{2}|+1}-\mathbb{E}[\ell_{\hat{\lambda}}(X_{n+1},Y_{n+1})\mid\hat{\lambda}]=\frac{B-\mathbb{E}[\ell_{\hat{\lambda}}(X_{n+1},Y_{n+1})\mid\hat{\lambda}]}{|\mathcal{I}_{2}|+1}\geq 0.$
By the tower property of expectation, there is
$\mathbb{E}[\tilde{\alpha}_{\ell}(\hat{\lambda})]\geq\mathbb{E}[\ell_{\hat{\lambda}}(X_{n+1},Y_{n+1})].$
[…]
Appendix F Additional Experiment Details
@@
We focus on four representative optimization paradigms for decision-making problems: ($i$) linear programming; ($ii$) the newsvendor problem (single-item inventory); (iiiiii) portfolio optimization (long-only, mean-CVaR proxy); and ($iv$) the shortest path problem (equivalently, a unit flow problem). The synthetic data used in our experiments focuses on settings where $X$ is degenerated, so we only have observations of $Y$. As we’ve mentioned before, this simplification does not limit the generality of our experimental findings, since both the Pareto frontier and its associated performance metrics are inherently unconditioned on $X$. The implementation is developed in Python, with core dependencies on the cvxpy (Diamond and Boyd, 2016) and cvxpylayers (Agrawal et al., 2019) libraries, which are used to solve convex optimization problems. We use the default parameter settings in their libraries across all experiments. Across all settings, the upper bound $B$ is heuristically set to the maximum observed regret for a large number ($1\times 10^{2}$) of simulations. $\lambda$ is varied from zero to the support radius of $Y$.
F.1 Experimental Setups
@@
In this section, we present the detailed experimental setups for the four optimization problems considered in our study. Across all settings, we specify the uncertainty set as $\mathcal{U}_{\lambda}=\{\tilde{y}:\|\tilde{y}-\mu\|_{\infty}\leq\lambda\}$, where $\mu$ denotes the mean of the distribution of the outcome variable $Y$. These optimizations are solved via convex solvers in our implementation.
@@
$\min_{z\in\mathbb{R}^{2}}\max_{y\in\mathcal{U}_{\lambda}}y^{\top}z\quad\text{s.t.}\quad Az\leq b.$
@@
$\min_{z\in\mathbb{R}_{+}}\max_{y\in\mathcal{U}_{\lambda}}\big[-p\min(y,z)+cz-v(z-y)^{+}\big].$
@@
$\min_{z\in\mathbb{R}_{+}^{m}}\max_{y\in\mathcal{U}_{\lambda}}y^{\top}z\quad\text{s.t.}\quad Az=b.\quad\Longleftrightarrow\quad\min_{z\in\mathbb{R}_{+}^{n_{\text{edge}}}}\mu^{\top}z+\lambda\,(w^{\top}z)\quad\text{s.t.}\quad Az=b,$
@@
F.2 Additional Experiment Results